10. (16) Find each indefinite integral using u-substitution: a. (x*(1-2x°)* dx b. fxcos (x2 - 1) dx
Answer #1

(x² (1-2x 8 Yde let 1-2x3 =u dCi) -2 d(x) - du 2(3x²dx) = du x²dx = du Irans Gerde) = sur dog tc [4] = 1 (1-223) + c 30 2² (1